# A Conflicts-free, Speed-lossless KAN-based Reinforcement Learning Decision System for Interactive Driving in Roundabouts

### v1

@@ III-B Structure of KAN
The core of the KAN architecture consists in using spline-based activation functions of the form:
@@
where $x_{i}$ is the input to the $i$-th neuron, $\text{spline}(x_{i};\theta_{i})$ represents the spline function parameterized by coefficients $\theta_{i}$, $b(x_{i})=\text{SiLU}(x_{i})=x_{i}/(1+e^{-x_{i}})$ is an activation function, and $\alpha_{i}$ and $\beta_{i}$ are learnable coefficients. Spline functions are piecewise polynomial functions with high expressiveness and can approximate any continuous function. By adjusting the parameters and coefficients of the spline function, KAN can approximate various complex nonlinear functions.
The coefficients of the spline, $\theta_{i}$, are updated through gradient descent methods using the loss function $\mathcal{L}(\theta)$ in (3). The update law is
@@
where $\eta$ is the learning rate. KAN also employs regularization strategies to mitigate the risk of overfitting:
@@
where $\mathcal{R}(\theta)$ is the regularization term added to $\mathcal{L}(\theta)$, which can control the complexity of the model whilst fitting the data and improve the generalization ability of the model. $\lambda_{1}$ and $\lambda_{2}$ are regularization coefficients. The $L_{1}$ regularization term, $\lambda_{1}\sum_{i}|\theta_{i}|$, promotes sparsity in the parameter matrix. Another $L_{1}$ regularization term, $\lambda_{2}\sum_{i}\sum_{j\neq i}|\theta_{i}-\theta_{j}|$, ensures smoothness in the parameter values across different neurons, facilitating model stability and preventing drastic changes in the output for minor fluctuations in the input data.
KAN also adopts parameter sharing among neurons defined as follows:
@@
where group includes indices of neurons sharing parameters, and $N_{\text{group}}$ is the number of neurons in this group. The shared parameters, $\theta_{\text{shared}}$, reduce the overall model complexity by averaging the parameters of neurons within the same group, enhancing computational efficiency and potentially improving the model’s generalization capabilities over similar features.
@@ III-C KAN Enhanced DQN
The integration of KAN within DQN, termed K-DQN, enables the network to better approximate the Q-function, especially in scenarios with complex action dynamics and reward functions. This capability translates into more robust learning and improved policy development in DRL tasks. To demonstrate why using the KAN network combined with DQN is superior to combining it with other reinforcement learning frameworks, it is essential to analyze the core differences in how these algorithms handle environments and tasks, as well as the characteristics of KAN activation functions.
To formulate the roundabout driving problem as a Markov Decision Process, we define the following key components:
State Space: The state space $\mathcal{S}$ consists of the ego vehicle’s position, velocity, and heading, as well as the relative positions, velocities, and headings of the surrounding vehicles within a certain range. The state at time $t$ is represented as
@@
where $p_{\textrm{EV}}(t)$, $v_{\textrm{EV}}(t)$, and $h_{\textrm{EV}}(t)$ denote the position, velocity, and heading of the ego vehicle at time $t$, and $p_{\textrm{NV}}^{i}(t)$, $v_{\textrm{NV}}^{i}(t)$, and $h_{\textrm{NV}}^{i}(t)$ represent the position, velocity, and heading of the $i$-th neighboring vehicle at time $t$.
Action Space: The action space $\mathcal{A}$ is discrete and consists of five high-level actions: faster, slower, idle, turn right, and turn left.
Reward Function: The reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is designed to encourage the ego vehicle to drive safely and efficiently through the roundabout. The reward function is defined as a weighted sum of several components:
@@
where $r_{\textrm{c}}$ penalizes collisions, $r_{\textrm{s}}$ rewards driving at high speeds, $r_{\textrm{l\_c}}$ penalizes excessive lane changes, $r_{\textrm{h}}$ encourages maintaining a safe distance from surrounding vehicles, and $r_{\textrm{a}}$ rewards reaching the target exit. The weights $w_{1},\ldots,w_{5}$ balance the importance of each component.
@@
where $W$ and $b$ are the weights and bias of the network, $f(x)$ is the output from the KAN layer, $j$ is the index of the output layer neuron, and $n$ is the total number of output layer neurons.
@@
[…]
@@
The optimal policy $\pi^{*}$ is derived from the optimal action-value function $Q^{*}$ by selecting the action that maximizes $Q^{*}$ for each state:
$\pi^{*}(a\mid s)=\arg\max_{a}Q^{*}(s,a).$ (24)
Therefore, K-DQN can approximate the optimal action-value function $Q^{*}(s,a)$ by optimizing over the loss function (3), and thus learn the optimal policy $\pi^{*}$. This demonstrates the effectiveness and superiority of KAN in enhancing DQN. The direct optimization approach of K-DQN, supported by the approximation theory and the quantitative error bound, enables K-DQN to better learn and optimize policies.
@@ IV Routes Planner and Actions Inspector
This section introduces the mechanisms to ensure the safety and efficiency during the interactive driving in the roundabout. Three subsections are included: the driving rules of HDVs, the action inspector, and the route planner.
@@ IV-A Driving Rules of HDVs
This subsection outlines the priority rules for HDVs operating within roundabouts, designed to maintain traffic flow and enhance safety. The rules are formulated to address typical scenarios encountered in roundabouts.
@@ IV-A1 Entry Rule
When an HDV approaches a roundabout entrance, it must yield to any vehicle already passing through the entrance it intends to use. This rule ensures that vehicles inside the roundabout maintain a smooth flow and reduces potential entry conflicts.
$\text{HDV}_{\text{entering}}\not\leftarrow\text{if }\exists\,\text{HDV}_{\text{passing}}$ (25)
Once inside the roundabout, ego HDVs (EHDV) are required to adjust their speeds according to the Intelligent Driver Model (IDM) to maintain a safe distance from their immediate front HDV (FHDV) until they reach their intended exit. This adjustment is crucial for preventing rear-end collisions and ensuring steady traffic flow within the roundabout. The IDM following rule is formulated as
$a_{\textrm{IDM}}=a_{\textrm{max}}[1-(v_{\textrm{FHDV}}/v_{e})^{4}-(h^{*}/h)^{2}]$ (26)
where $a_{\textrm{max}}$ is the maximum acceleration of EHDV, $v_{\textrm{FHDV}}$ is the velocity of FHDV, $v_{e}$ is the expected velocity of EHDV, and $h$ is the real gap between EHDV and FHDV. $h^{*}$ is the desired gap between EHDV and FHDV with the formula
$h^{*}=h_{e}+v_{\textrm{AV}}T_{e}-\frac{v_{\textrm{AV}}{\Delta v}}{2\sqrt{a_{\textrm{max}}c}}$ (27)
where $h_{e}$ is the expected space with FHDV, $v_{\textrm{AV}}$ is the speed of the AV, $T_{e}$ is the expected time gap, $\Delta v$ is the velocity difference between EHDV and FHDV, and $c$ is the comfortable deceleration.
@@ IV-B Route Planner
The integrated route planner for the EV comprises initial-lane selection decisions, a path-planning algorithm, and a lane-change selection mechanism. The initial-lane selection is guided by the TTC metric for each lane, ensuring safety and efficiency from the start. The path planning algorithm employs a node-based shortest path calculation to determine the most optimal route. The lane-change selection mechanism is driven by a proposed lane change cost formula, facilitating effective and strategic lane changes.
@@ IV-B1 Initial-Lane Selection
$\text{TTC}_{\text{inner}}=\frac{\text{Distance to HDV}_{\text{inner}}}{\text{Speed of EV}-\text{Speed of HDV}_{\text{inner}}},$
$\text{TTC}_{\text{outer}}=\frac{\text{Distance to HDV}_{\text{outer}}}{\text{Speed of EV}-\text{Speed of HDV}_{\text{outer}}}.$ (28)
@@
When no HDVs are present in the roundabout, the AV will directly select the inner lane due to its shorter radius, which reduces the distance from the entrance to the outlet.
$\bullet$ One HDV in outer lane: Fig. 4(a) illustrates this scenario. In this scenario, the EV will compute the TTC to maintain a safe distance from surrounding vehicles and merge into the inner lane.
@@
$\text{Lane}_{\text{selected}}=\!\begin{cases}\text{Inner lane}&\!\!\text{if }\text{TTC}_{\text{inner}}\geq\text{TTC}_{\text{outer}}\\ \text{Outer lane}&\!\!\text{otherwise}\end{cases}\!.\!\!$ (30)
@@
$\bullet$ Multiple HDVs in both lanes: This scenario is most complex, as both inner-lane and outer-lane have two HDVs. When multiple vehicles (in this paper more than 2 vehicles are defined as multiple case) are present in both lanes, a weighted decision based on TTC and Total Driving Time (TDT) to the outlet is used:
$\text{TTC}_{\text{weighted}}=w_{1}\cdot\text{TTC}_{\text{nearest}}+w_{2}\cdot\text{TDT},$
@@
The weights $w_{1}$ and $w_{2}$ are predetermined based on the traffic model preferences. The lane with the lower weighted score is selected to maximize safety and avoid the lane with longer total travel time of HDVs, which prevents the forward driving of AV. The entire initial-lane selection process is summarized in the Algorithm 1.
(a) Scenario with one HDV in the outer lane
(b) Scenario with one HDV in both lanes
Figure 4: Lane selection cases. (a) shows the case with one HDV in the outer lane, requiring the EV to select the safest lane by the higher TTC; (b) shows the case with one HDV in both lanes, and the EV selects the lane by the higher TTC.
Algorithm 1 Action Priority List for the EV
1:$L$: The number index of lanes in the roundabout. $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$: Coefficients for computing the priority scores based on the current situation. $T_{n}$: The number of prediction steps for trajectory planning. $A_{t,i}$: Set of feasible actions for vehicle $i$ at time step $t$.
2:$\mathit{P_{t}}$: Priority list of actions for the EV.
@@ IV-B2 In-Roundabout Lane Selection
After entering the roundabout and selecting an initial lane, the next stage is path planning. We adopt a modified Breadth-First Search (BFS) [30] method that considers both distance and traffic conditions to compute the optimal path from a start point to a target within a graph structure, where nodes represent intersections in the road network, and edges represent drivable roads. The modified BFS algorithm uses a cost function that takes into account both the distance and traffic density:
@@ IV-B3 Lane Selection Mechanism
As the scenario illustrated in Fig. 4(b), traffic density ($\mathcal{D}$) and lane-change cost ($\mathcal{C}$), are computed. The computation of two parameters are as follows:
@@
$\mathbf{1}_{\text{condition}}=\begin{cases}1&\text{if the condition is true}\\ 0&\text{otherwise}\end{cases}.$ (37)
@@
where $D(EV,NV)$ is the distance between EV and NV, $p_{\textrm{EV}}(t)$ is EV’s position at time step $t$, and $p_{\textrm{NV}}(t)$ is NV’s position at time step $t$.
Based on the above parameters, the lane is selected using
@@
If the costs for both lanes are equal, the decision is further optimized based on the vehicle’s current position. This approach enhances the efficiency and safety of vehicle control around a roundabout by integrating real-time traffic conditions with the potential risks of lane changes. By utilizing the modified BFS algorithm, the algorithm optimizes the selection of paths. Besides, the route planner employs an edge selection function, creating a comprehensive solution for autonomous vehicle navigation in roundabouts.
IV-C Action Inspector
Each ego vehicle (EV) plans its acceleration through the roundabout and it is encouraged to accelerate during driving. However, the collision risk should be measured. To this end, the EV would predict the trajectories of NVs when the distance between the EV and any NV is less than the $(D_{safe})$ for the EV to change lane safely as depicted in Fig. 4 (b).
If there are overlaps between EV’s future trajectories and NVs’ trajectories, the EV will use IDM (26) to follow the nearest preceding vehicles, and to synchronize with the overall traffic flow. The following mode would generate a penalty if the EV’s velocity is below the expected velocity for more than three seconds. This mechanism enables the EV to avoid conflicts and understand the requirement to drive to the outlet as quickly as possible.
IV-C1 Safety Margin Calculation
This margin is used to guide decision-making when selecting driving actions. As vehicles maintain a wider angle relative to each other while in proximity, the likelihood of their paths intersecting decreases. Therefore, the safety margin for each vehicle’s maneuver is defined as the minimum difference in relative angle, $D_{a}$, or the shortest time until a potential collision could occur.
Safety $\text{Safety Margin}=\min\limits_{a\in A_{\text{feasible}}}D_{\text{a},a,k}$ (40)
IV-C2 Decision-Making Criteria
For each decision point, the AV will calculate safety margins for multiple options. If the safety margins are equivalent, the AV will prefer the lane that optimizes the route, typically the inner lane in roundabouts due to the shorter path length to the exit.
IV-C3 Dangerous Action Replacement
When the action inspector detects a dangerous action that may lead to a collision, it replaces the action with the next highest priority action from the list $P_{t}$. This process ensures that the EV always selects the safest possible action while still being efficient. If no safe action is found in $P_{t}$, the EV will use IDM to follow the nearest preceding vehicle until a safe action becomes available.
Algorithm 2 Action execution for EV in a Roundabout
1:$\mathit{P_{t}}$: Priority list of actions for the EV, initialized and populated as per the previous algorithm.
2:$a_{t,i}$: The optimal action for vehicle $i$ at time step $t$, chosen and executed from the priority list.
@@ IV-C3 Dangerous Action Replacement
5:     for $NV$ in $N_{V}$ and $D(EV,NV)\leq D_{safe}$ do
6:         if $EV_{a_{t}}$ and $NV$ trajectories overlap in $T_{n}$ then
7:              if $NV$ in same lane and in front then
@@
9:              else if $NV$ in adjacent lane then
@@
IV-C4 Update Rule

### v2

@@ III-B Structure of KAN
KAN’s core uses spline-based activation functions of the form:
@@
where $x_{i}$ is the input to the $i$-th neuron, $\text{spline}(x_{i};\theta_{i})$ represents the spline function parameterized by coefficient $\theta_{i}$, $b(x_{i})=\text{SiLU}(x_{i})=x_{i}/(1+e^{-x_{i}})$ is an activation function, and $\alpha_{i}$ and $\beta_{i}$ are learnable coefficients. Spline functions are piecewise polynomials that can approximate any continuous function. By tuning their parameters, KAN can model complex nonlinear functions.
The coefficient $\theta_{i}$ is updated via gradient descent on the loss $\mathcal{L}(\theta)$ in (3), following the update rule:
@@
with the learning rate $\eta$.
KAN uses regularization to reduce overfitting by adding to $\mathcal{L}(\theta)$ the term:
@@
where $\lambda_{1}$ and $\lambda_{2}$ are regularization coefficients. The $L_{1}$ term $\lambda_{1}\sum_{i}|\theta_{i}|$ promotes sparsity, while $\lambda_{2}\sum_{i}\sum_{j\neq i}|\theta_{i}-\theta_{j}|$ enforces smoothness across neurons, enhancing stability. Overall, adding $\mathcal{R}(\theta)$ controls model complexity and improves generalization.
KAN also uses parameter sharing among neurons, defined as:
@@
where group indexes neurons sharing parameters, and $N_{\text{group}}$ is the group size. Shared parameters $\theta_{\text{shared}}$ average neuron parameters, reducing model complexity, improving efficiency, and enhancing generalization.
@@ III-C KAN Enhanced DQN
Integrating KAN into DQN (K-DQN) enhances Q-function approximation, boosting learning robustness and policy performance in complex DRL tasks. To justify pairing KAN with DQN over other RL methods, it’s crucial to analyze their differences in handling environments and KAN’s activation function traits.
To model roundabout driving as a Markov Decision Process, we define these key components:
State Space $\mathcal{S}$: It consists of the ego vehicle’s position, velocity, and heading, as well as the relative positions, velocities, and headings of the surrounding vehicles within a certain range. The state at time $t$ is represented as
@@
where $p_{\textrm{EV}}(t)$, $v_{\textrm{EV}}(t)$, and $h_{\textrm{EV}}(t)$ denote the position, velocity, and heading of the ego vehicle, while $p_{\textrm{NV}}^{i}(t)$, $v_{\textrm{NV}}^{i}(t)$, and $h_{\textrm{NV}}^{i}(t)$ are the position, velocity, and heading of the $i$-th neighboring vehicle.
Action Space $\mathcal{A}$: It is discrete and consists of five high-level actions: faster, slower, idle, turn right, and turn left.
Reward Function $r$: It encourages the ego vehicle to drive safely and efficiently through the roundabout, designed as:
@@
where $r_{\textrm{c}}$ assigns a large penalty (-100) for collisions, $r_{\textrm{s}}$ provides continuous feedback proportional to the vehicle’s speed ($v/v_{\text{max}}$), $r_{\textrm{l\_c}}$ adds a small negative value (-10) for each lane change to prevent unnecessary maneuvers, $r_{\textrm{h}}$ encourages maintaining safe distances by scaling with inverse headway time, and $r_{\textrm{a}}$ gives a positive reward (+200) for reaching the target exit. The weights are empirically set as $w_{1}=1.0$, $w_{2}=0.3$, $w_{3}=0.2$, $w_{4}=0.3$, and $w_{5}=0.2$ to balance safety with efficiency.
@@
where $W$ and $b$ are the weight and bias of the network, $f(x)$ is the output from the KAN layer, $j$ is the index of the output layer neuron, and $n$ is the total number of output layer neurons.
@@
where $C$ is a constant and $G$ is the grid size. The magnitude of derivatives up to order $m$ is measured by the $C^{m}\text{-norm}$ as
@@
The optimal policy $\pi^{*}$ selects actions maximizing the optimal Q-value $Q^{*}$ for each state:
$\pi^{*}(a\mid s):=\arg\max_{a}Q^{*}(s,a).$ (24)
Thus, K-DQN can approximate $Q^{*}(s,a)$ by minimizing the loss function (3), enabling it to learn the optimal policy $\pi^{*}$. This highlights KAN’s effectiveness in enhancing DQN through direct optimization and strong theoretical guarantees.
III-D Computational Complexity Analysis
Integrating KAN into DQN adds computational overhead from spline-based activations. While traditional DQN’s forward pass has complexity $O(LN^{2})$ for $L$ layers and $N$ neurons, K-DQN’s complexity increases to $O(LN^{2}+LNS)$, where $S$ is the number of spline segments.
During training, the backward propagation in traditional DQN has complexity $O(LN^{2})$. For K-DQN, the gradient computation through spline functions adds an overhead, resulting in $O(LN^{2}+LNS)$ complexity. However, two factors help mitigate this computational cost: 1) Parameter sharing among neurons reduces the number of parameters to be updated, lowering the practical computational burden to approximately $O(LN^{2}+LN\bar{S})$, where $\bar{S}<S$ is the effective number of unique spline segments. 2) The improved approximation capabilities of KAN typically require fewer training iterations for convergence. Our empirical results show that K-DQN achieves equivalent performance with approximately 30% fewer training iterations compared to traditional DQN.
During inference, with fixed spline coefficients, forward pass complexity reduces to $O(LN^{2}+LN)$, causing only a slight increase in decision time over traditional DQN.
@@ IV Routes Planner and Actions Inspector
This section presents mechanisms for safe and efficient roundabout driving, covering HDV driving rules, the action inspector, and the route planner.
@@ IV-A Driving Rules of HDVs
This subsection outlines HDV priority rules in roundabouts to maintain traffic flow and safety, addressing common scenarios.
@@ IV-A1 Entry Rule
When an HDV nears a roundabout, it must yield to vehicles already in the entrance it plans to use, ensuring smooth traffic flow and minimizing conflicts. Such a rule is described as:
$\text{HDV}_{\text{entering}}\not\leftarrow\text{if }\exists\,\text{HDV}_{\text{passing}}.$ (25)
Inside the roundabout, ego HDVs (EHDVs) must adjust their speed per the Intelligent Driver Model (IDM) policy in (26) to maintain a safe gap from the front HDV (FHDV) until exiting, preventing rear-end collisions and ensuring smooth flow.
$a_{\textrm{IDM}}=a_{\max}[1-(v_{\textrm{FHDV}}/v_{e})^{4}-(h^{*}/h)^{2}]$ (26)
where $a_{\textrm{max}}$ is the maximum acceleration of EHDV, $v_{\textrm{FHDV}}$ is the velocity of FHDV with the desired value $v_{e}$, and $h$ is the real gap between EHDV and FHDV. $h^{*}$ is the desired gap between EHDV and FHDV with the formula
$h^{*}=h_{e}+v_{\textrm{AV}}T_{e}-v_{\textrm{AV}}{\Delta v}/(2\sqrt{a_{\max}c})$ (27)
where $h_{e}$ is the expected space to FHDV, $v_{\textrm{AV}}$ is AV’s speed, $T_{e}$ is the desired time gap, $\Delta v$ is the velocity difference between EHDV and FHDV, and $c$ is the comfortable deceleration.
@@ IV-B Route Planner
The integrated route planner for the ego vehicle (EV) comprises initial-lane selection decisions, a path-planning algorithm, and a lane-change selection mechanism. The initial-lane selection is guided by the TTC metric for each lane, ensuring safety and efficiency from the start. The path planning algorithm employs a node-based shortest path calculation to determine the most optimal route. The lane-change selection mechanism is driven by a proposed lane change cost formula, facilitating effective and strategic lane changes.
@@ IV-B1 Initial-Lane Selection
$\text{TTC}_{\text{inner}}=\frac{\text{Distance to HDV}_{\text{inner}}}{\text{Speed of EV}-\text{Speed of HDV}_{\text{inner}}},$ (28)
$\text{TTC}_{\text{outer}}=\frac{\text{Distance to HDV}_{\text{outer}}}{\text{Speed of EV}-\text{Speed of HDV}_{\text{outer}}}.$
@@
With no HDVs present, the AV selects the inner lane for its shorter, more direct path through the roundabout.
$\bullet$ One HDV in outer lane: Fig. 4(a) illustrates this scenario, where the EV computes the TTC to maintain a safe distance from surrounding vehicles and merge into the inner lane.
@@
$\text{Lane}_{\text{selected}}=\!\begin{cases}\text{Inner lane},&\!\!\text{if }\text{TTC}_{\text{inner}}\geq\text{TTC}_{\text{outer}}\\ \text{Outer lane},&\!\!\text{otherwise}\end{cases}\!.\!\!$ (30)
@@
$\bullet$ Multiple HDVs in both lanes: This is the most complex scenario, with two HDVs in both lanes. When multiple vehicles (more than two) are present, a weighted decision based on TTC and Total Driving Time (TDT) to the outlet is applied:
$\text{TTC}_{\text{weighted}}=w_{1}\cdot\text{TTC}_{\text{nearest}}+w_{2}\cdot\text{TDT},$ (31)
$\text{TDT}=\sum\text{Driving time of each HDV to EV's outlet},$
where $w_{1}$ and $w_{2}$ are predefined weights reflecting traffic model preferences. The lane with the lower score is chosen to enhance safety and avoid HDV delays. The full selection process is outlined in Algorithm 1.
Figure 4: Lane selection cases: (a) One HDV in the outer lane—EV chooses the safer lane with higher TTC; (b) One HDV in each lane—EV selects the lane with higher TTC.
Algorithm 1 Action priority list for EV
1:$L$: Lane index in the roundabout; $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$: Coefficients for priority score computation; $T_{n}$: Prediction horizon for trajectory planning; $A_{t,i}$: Feasible actions for vehicle $i$ at time $t$
@@ IV-B2 In-Roundabout Lane Selection
After entering the roundabout and selecting an initial lane, the next stage is path planning. We adopt a modified Breadth-First Search (BFS) [35] method that considers both distance and traffic conditions to compute the optimal path from a start point to a target within a graph structure, where nodes represent intersections in the road network, and edges represent drivable roads. The modified BFS algorithm uses the cost function:
@@ IV-B3 Lane Selection Mechanism
As the scenario illustrated in Fig. 4(b), traffic density ($\mathcal{D}$) and lane-change cost ($\mathcal{C}$), are computed. Additionally, to align with real-world driving behaviors where vehicles preparing to exit typically move to the outer lane in advance, we implement a dynamic priority adjustment mechanism.
@@
$\mathbf{1}_{\text{condition}}=\begin{cases}1,&\text{if the condition is true}\\ 0,&\text{otherwise}\end{cases}.$ (37)
@@
where $D(EV,NV)=\|p_{\textrm{EV}}(t)-p_{\textrm{NV}}(t)\|_{2}$ is the distance between EV and non-ego vehicle (NV), $p_{\textrm{EV}}(t)$ is EV’s position, and $p_{\textrm{NV}}(t)$ is NV’s position. This inter-vehicle distance-based cost calculation adapts naturally to different lane geometries without parameter adjustment.
$\bullet$ Outer Lane Preference $\omega(d)$ is designed to encourage timely transitions to the outer lane when approaching the target exit. This preference weight is defined as:
$\omega(d)=\begin{cases}0,&\text{if }d<0.5d_{\text{total}}\\ \beta\cdot\frac{d-0.5d_{\text{total}}}{0.5d_{\text{total}}},&\text{otherwise}\end{cases}$ (39)
where $d$ is the distance traveled from the entrance, $d_{\text{total}}$ is the total distance to the target exit, and $\beta$ is a weighting parameter (set to 0.3 in our experiments) that controls the strength of outer lane preference. The lane choice $l_{c}$ is defined as:
$l_{c}:=\arg\min_{l\in\{0,1\}}\left(\mathcal{D}(n,l)+\mathcal{C}(n,l)-\omega(d)\cdot\mathbf{1}_{(l=\text{outer})}\right).$ (40)
This formulation ensures that as the vehicle approaches its exit (when $d>0.5d_{\text{total}}$), the outer lane becomes increasingly preferable, reflecting realistic driving behavior while maintaining safety through density and lane change cost terms. When lane costs are equal, the decision is refined based on the vehicle’s position. This enhances both safety and efficiency by integrating real-time traffic conditions with potential lane change risks, while preserving natural driving tendencies. The route planner leverages a modified BFS algorithm with an edge selection function to identify optimal paths, providing a robust solution for autonomous navigation in roundabouts.
Algorithm 2 Action execution for EV in roundabout
@@
[…]
@@
5:  for NVNV in $N_{V}$ and $D(EV,NV)\leq D_{safe}$ do
6:   if $EV_{a_{t}}$ and NVNV trajectories overlap in $T_{n}$ then
7:     if NVNV in same lane and in front then
@@
9:     else if NVNV in adjacent lane then
@@
IV-C Action Inspector
Each EV plans its acceleration through the roundabout, with an incentive to accelerate for efficiency. However, to ensure safety, the EV predicts the trajectories of nearby NVs whenever their distances fall below a safety threshold $D_{safe}$, as shown in Fig. 4(b). This safety distance accounts for vehicle dimensions by expanding the EV’s boundary by half of its width ($1.05\,\mathrm{m}$) and length ($2.35\,\mathrm{m}$), ensuring sufficient clearance for maneuvers. The vehicle-dimension-based safety calculation ensures robustness to varying lane widths. If predicted trajectories overlap, the EV switches to a following mode using the IDM policy (26), adapting to the nearest preceding vehicle and synchronizing with traffic flow. A penalty is imposed if the EV’s speed remains below the expected value for more than three seconds, encouraging timely progression while maintaining safety.
1) Safety Margin Calculation: This margin is used to guide decision-making when selecting driving actions. As vehicles maintain a wider angle relative to each other while in proximity, the likelihood of their paths intersecting decreases. Therefore, the safety margin for each vehicle’s maneuver is defined as the minimum difference in relative angle, $D_{a}$, or the shortest time until a potential collision could occur.
$\text{Safety Margin}=\min\limits_{a\in A_{\text{feasible}}}D_{\text{a},a,k}$ (41)
2) Decision-Making Criteria: For each decision point, the AV will calculate safety margins for multiple options. If the safety margins are equivalent, the AV will prefer the lane that optimizes the route, typically the inner lane in roundabouts due to the shorter path length to the exit.
3) Dangerous Action Replacement: When a dangerous action is detected, the inspector replaces it with the next highest-priority action from $P_{t}$, ensuring the EV chooses the safest option. If none are safe, the EV follows the nearest vehicle using IDM until a safe action appears.